a soccer field is a rectangle 90 meters wide and 120 meters long the coach ask the players to run from one corner to the corner diagonally across the field how far do the player run? do you know because i missed when my teacher was teaching this now im a day behind
Answer by JulietG(1812)

A diagonal through a rectangle is the hypotenuse of a right angle triangle.
We know that a^2 + b^2 = c^2 for a right triangle where a and b are the vertical and horizontal measurements.(the ^2 means "squared")
You're given a as 120m.
You're given b as 90m
Find c.
120^2 + 90^2 = c^2
14400 + 8100 = c^2
22500 = c^2
Square root of each side:
150 = c (the diagonal)